OpenStudy (anonymous):

giving medal for best answer Solve 7 and one-ninth minus 4 and two-thirds.

OpenStudy (anonymous):

try it urself first

OpenStudy (anonymous):

Ill help you now..

OpenStudy (anonymous):

(7 + 1/9) - (4 + 2/3) = (7(9/9) + 1/9) - (4(3/3 + 2/3)) = (63/9 + 1/9) - (12/3 + 2/3) = 64/9 - 14/3 = 64/9 - (3/3)(14/3) = 64/9 - 42/9 = 22/9 = (18 + 4) / 9 = 18/9 + 4/9 = 2 + 4/9 Ans: Two and four ninths.

OpenStudy (anonymous):

Ty (:

OpenStudy (anonymous):

well done :)

OpenStudy (anonymous):

@chandanjha ty